# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

### v1

@@ 3.1 Randomized Matrix Sketching
    where $\bm{\mathsf{D}}\in\mathbb{F}^{m\times m}$ is a random diagonal matrix populated with i.i.d. Rademacher random variables, $\bm{\mathsf{F}}\in\mathbb{F}^{m\times m}$ is a discrete Fourier or Walsh-Hadamard transform, and $\bm{\mathsf{S}}\in\mathbb{F}^{d\times m}$ is the transpose of a SparseCol matrix with exactly $\zeta$ nonzero entries per column as in the SparseCol matrix above.
These distributions yield randomized DRMs that work well in practice, with theoretical guarantees summarized in Table 1. Gaussian test matrices are oblivious subspace embeddings that enjoy a superior lower bound on the requisite embedding dimension $d$ to guarantee (21), but the costs of storage and matrix-vector products for Gaussians are much more expensive than for “structured” embeddings like SRTT or sparse sign matrices. Put simply, Gaussian sketching makes for straightforward analysis, but randomized sketching with structured random matrices makes for better performance. The reader is referred to (martinsson2020, Sections 8,9) for detailed discussion.
@@
Table 1: Summary of coordinate-mixing randomized DRMs commonly used for (row-)sketching rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace embeddings (OSE). Theoretical lower bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy (21) are provided, as well as asymptotic complexities to store $\bm{\mathsf{\Gamma}}$ and to apply $\bm{\mathsf{\Gamma}}$ to a vector $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptm}$, cf. Camano2025_OSI; martinsson2020; woodruff2015.
While the embedding dimensions in Table 1 guarantee the oblivious subspace embedding (OSE) property (21), it has been observed that smaller values of $d$ are often sufficient for many linear algebra tasks in practice when utilizing any of the randomized DRMs of Table 1 (with the exception of CountSketch), e.g., $d=k+10$, cf. Sect. 4.1. Recently, this observation was thoroughly investigated from the perspective of oblivious subspace injections, versus embeddings, which relaxes the OSE property to the following weaker one.
A random matrix $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an oblivious subspace injection (OSI) of $\bm{\mathsf{A}}$ with injectivity $\alpha$ if it satisfies the injectivity property
@@
With the OSI property guiding our choice of random sketching matrix, we enjoy more favorable sketching dimensions for the random matrices considered above. In the OSI framework, Gaussian random matrices are still the gold standard for approximation accuracy, but suffer the same performance drawbacks in terms of storage and matrix-vector product, or mat-vec, costs. Structured random matrices offer significantly better performance, and specifically, sparse random matrices exhibit the best empirical performance, as demonstrated by many practical investigations Chen2025_sparseGPU; ChenakkodDereZinskiDongRudelson2024_OptimalEmbeddingSparse; EpperlyMeierNakatsukasa2024_FastRandomizedLeastSquares; halko2011; KaneNelson2014; martinsson2020; nelson2013; Tropp2015_MatrixConcentration; woodruff2015, and now also supported by the theoretical guarantees of Camano2025_OSI; Tropp2025_OSI summarized in Table 2.
In general, sparse random matrices are recommended for sketching a dense unstructured input matrix, due to the reduced storage costs and significant acceleration afforded by high-performance sparse arithmetic libraries. To that end, sketching with SparseRTT matrices, or sparsified SRTTs, is more computationally efficient than with standard SRTTs or Gaussians. Compared to the SRTT, the SparseRTT matrix can reduce data movement and better facilitate parallelism at the hardware level. However, the SparseStack matrix is typically faster and thus recommended over SparseRTTs whenever sparse linear algebra packages are available.
As an additional boon, the OSI values of $d$ that are reported in Table 2 are even still too pessimistic for practical implementation purposes. In particular, the results of Camano2025_OSI provide compelling evidence that SparseStack matrices are the superior choice of DRM in practice; it is conjectured that SparseStack matrices can achieve the $\frac{1}{2}$-OSI property with $d=O(k)$ and constant sparsity $\zeta=O(1)$. This conjecture is supported numerically, with $d=2k$ and $\zeta=4$ used in the experiments of Camano2025_OSI to demonstrate that the low-rank approximations obtained from the SparseStack sketches are of comparable accuracy to those obtained from Gaussian sketches.
Finally, we note that for input matrices that arise from tensors (cf. Sect. 5.3), we can construct Khatri-Rao sketching matrices that exploit underlying Kronecker product structure Camano2025_OSI. As we show in Sect. 6.1, these sketching matrices are formed as Kronecker products of isotropic random vectors. Often, these random vectors are drawn from (real) Gaussian and Rademacher distributions, but it is advised in Camano2025_OSI that the complex spherical distribution be used to avoid near-zero injectivity $\alpha$, which can occur in worst-case scenarios. For our sketching matrix constructions in Sect. 6, we do not assume any specific isotropic distribution unless otherwise specified.
@@
Table 2: Summary of coordinate-mixing randomized DRMs commonly used for row-sketching rank-$k$ $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace injections (OSI). Theoretical bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSI property (25) vs. the OSE property (21) are provided, cf. Camano2025_OSI; Tropp2025_OSI. For OSIs, we enforce a constant injectivity parameter, i.e. $\alpha\geq 1/2$, which is generally not attainable for OSEs due to the dilation parameter. For the OSE column of SparseRTTs, the dagger $\dagger$ signifies that a different sparse matrix is used for the OSE property, cf. (Camano2025_OSI, Section 4.1).
@@ 3.2 Randomized Matrix Sampling
In the previous section, we observed that a random sketch $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ (or $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$) is a new matrix whose rows (or columns) are random linear combinations of the rows (or columns) of $\bm{\mathsf{A}}$. However, we can also construct randomized DRMs that sample coordinates of $\bm{\mathsf{A}}$ according to a probability distribution. We use the separate term randomized sampling to refer to these types of maps.
@@
  • Uniform ((martinsson2020, Section 9.6.2), (tropp2011b, Lemma 3.4), Kannan2017):
    $p_{i}=1/m.$
  • Squared (row) norms (Frieze2004):
    $p_{i}=\|\bm{\mathsf{a}}_{i}\|_{2}^{2}/\|\bm{\mathsf{A}}\|^{2}_{\text{F}},$
    where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$,
@@
  • Determinantal point processes (DPP) ((DerezinskiMahoney2021, Section 5.2), kulesza2012determinantal): If $S\subseteq[m]$, $|S|=d$,
@@
    where $(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}$ is the sub-matrix of $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\in\mathbb{F}^{m\times m}$ indexed by $S$ and all indices $S$ are selected at once. However, in practice, vectors (rows) corresponding to indices $S$ are selected iteratively, as in the algorithm presented in Hough_DPP_2005. Let $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}=\sum_{i}\lambda_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ denote the eigendecomposition with $\lambda_{i}\in[0,1]$ for all $i$. Let $s_{i}\sim\textup{Bernoulli}(\lambda_{i})$, and let $\bm{\mathsf{U}}$ be the matrix with columns $\bm{\mathsf{u}}_{i}$. By (Hough_DPP_2005, Theorem 7) and (DerezinskiMahoney2021, Theorem 12), we can use $\sum_{i}s_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ as a proxy for DPP sampling as follows, letting $\bm{\mathsf{v}}_{i}^{*}$ represent the $i$th row of $\bm{\mathsf{U}}$:
@@
  • Nuclear norm maximization (Altschuler2016GreedyCSS; Fornace2024; Ordozgoiti2018IterativeCSS; Ordozgoiti2019RegularizedCSS): While neither a traditional sketching or sampling method, we include nuclear norm maximization in this list because of its comparable performance to DPP sampling. Let $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$, $S\subseteq[m]$, $|S|=d$, and define
    $\mathcal{L}_{\bm{\mathsf{K}}}(S)=\textup{trace}\left(\bm{\mathsf{K}}_{:,S}(\bm{\mathsf{K}}_{S,S})^{-1}\bm{\mathsf{K}}_{S,:}\right)=\textup{trace}\left((\bm{\mathsf{K}}^{2})_{S,S}(\bm{\mathsf{K}}_{S,S})^{-1}\right).$
    Maximizing $\mathcal{L}_{\bm{\mathsf{K}}}(S)$ over all possible subsets of $[m]$ of size $d$ is infeasible as in the DPP setting, so we compute approximate nuclear scores using random sketches of $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$, and the identities
    $\textup{diag}(\bm{\mathsf{K}}^{2})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{Kx}})^{\odot 2}],\hskip 5.69054pt\textup{diag}(\bm{\mathsf{K}})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{\bm{\mathsf{A}}\bm{\mathsf{x}}}})^{\odot 2}],$
    where $\odot 2$ denotes element-wise squaring. Namely, we compute the averages
    $\textup{diag}(\bm{\mathsf{K}}^{2})/\textup{diag}(\bm{\mathsf{K}})\approx\frac{(\bm{\mathsf{K}}\bm{\mathsf{\Omega}})^{\odot 2}\bm{\mathsf{1}}}{(\bm{\mathsf{A}}\bm{\mathsf{\Phi}})^{\odot 2}\bm{\mathsf{1}}}\Rightarrow\bm{\mathsf{K}}^{2}_{\ell,\ell}/\bm{\mathsf{K}}_{\ell,\ell}\approx\frac{\|[\bm{\mathsf{K}}\bm{\mathsf{\Omega}}]_{\ell,:}\|^{2}_{2}}{\|[\bm{\mathsf{A}}\bm{\mathsf{\Phi}}]_{\ell,:}\|^{2}_{2}},$
    where $\bm{\mathsf{\Omega}},\bm{\mathsf{\Phi}}$ are Gaussian matrices with $O(\log m)$ columns. We then update $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$ by Schur complements with respect to the selected index, and repeat d1d-1 times.
@@
We can express the action of randomized sampling as a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$. The lower bounds on the minimal dimension required for a random sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) are summarized in Table 3, along with the computational complexity of computing the associated probabilities. Unlike the bounds in Table 1, the values of $d$ required for a randomized sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) depend not only on the rank $k$ of $\bm{\mathsf{A}}$ and the distortion parameter $\varepsilon$, but also, in some cases, on quantities associated with the matrix $\bm{\mathsf{A}}$ that are usually very expensive to compute.
@@
For uniform sampling, the requisite embedding dimension $d$ depends on the coherence $\mu$ of $\bm{\mathsf{A}}$, defined as $\mu(\bm{\mathsf{A}})=\max_{1\leq i\leq m}\frac{n}{m}l_{i}$, or the maximum leverage score $l_{i}$ in (26). The coherence is a measure of how evenly information in the matrix is distributed; high coherence indicates that its key information is very localized and that uniform sampling may perform poorly as a result; see, e.g., Ipsen14 or (martinsson2020, Section 9.6.2). In spite of this drawback, uniform sampling has become a popular choice for large-scale linear algebra computations as a cost-effective oblivious subspace embedding (OSE) method. However, it is unclear how to best choose the embedding dimension $d$ in practice since the coherence is typically unavailable, cf. (martinsson2020, Section 9.6.4).
@@
Squared norm sampling and leverage score sampling also yield randomized subspace embeddings, but not oblivious ones, though we can still analyze the embedding dimensions required for them to satisfy (21). For squared norm sampling, the value of $d$ depends on the condition number $\kappa$ of $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$, defined as the ratio of its largest to smallest eigenvalue, which is not readily available in practice. The embedding dimension for leverage score sampling only depends on $k$ and $\epsilon$, but this independence comes at great computational expense to compute the sampling probabilities. In particular, we note that leverage scores are as expensive to directly compute as $\texttt{svd}(\bm{\mathsf{A}})$, rendering their direct computations infeasible for most practical purposes. However, there exist fast methods to approximate them, cf. AlaouiMahoney15; Chen21; DrineasMagdonMahoneyWoodruff12; MuscoMusco17.
@@
DPP sampling is known to have nearly optimal theoretical guarantees, cf. derezinski2020improved. However, naive implementations incur a cost of $O(m^{3})$, which makes DPP sampling practically infeasible. In derezinski2019exact, the DPP-VFX algorithm is presented, which, given access to $\bm{\mathsf{A}}\bm{\mathsf{A}}^{\top}$, samples exactly from a DPP distribution with a pre-processing cost of $O(mk^{10}+k^{15})$. In particular, the DPP-VFX algorithm was the first to achieve a cost of $O(m\cdot\textup{poly}(k))$ for the first sample and $O(\textup{poly}(k))$ for subsequent samples; however, there still exists a runtime bottleneck of $\Omega(m)$ due to the computation of $m$ marginals for the associated sampling distribution. The same authors improved on DPP-VFX with the $\alpha$-DPP algorithm of calandriello2020sampling, which avoids the $\Omega(m)$ bottleneck by incorporating uniform sub-sampling before computing marginals.
While nuclear norm maximization is a greedy method of choosing columns based on posterior error minimization, the authors of Fornace2024 present a “matrix-free” algorithm for fast implementation with randomization. Its performance is comparable to that of DPP sampling, and nuclear norms could be adapted into a sampling distribution, so we include it in this section for side-by-side comparisons in Table 3. Namely, in Fornace2024, randomized sketching is used to compute the approximate scores that dictate column selection. We note that the requisite embedding dimension $d$ for nuclear norm maximization depends on the quantity $\nu=\frac{\textup{trace}(\bm{\mathsf{K}})-\textup{trace}^{(k)}(\bm{\mathsf{K}})}{\textup{trace}^{(k)}(\bm{\mathsf{K}})}$, where $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$ and $\textup{trace}^{(k)}(\bm{\mathsf{K}})$ is the sum of its $k$ largest eigenvalues. For the complexity, we also include the dependence on the quantity $\phi$, which is the maximum cost of a mat-vec with $\bm{\mathsf{K}}$ or $\bm{\mathsf{A}}$, cf. (Fornace2024, Corollary 4.1).
@@
Nuclear norm max $d=\Omega(k/\varepsilon+k-1)(\log(\nu^{-1})+\log(1/\varepsilon)))$ $O(md^{2}\log m+d\log m\psi)$
Table 3: Summary of randomized sampling distributions commonly used to sample rows of rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Theoretical lower bounds on the embedding dimension $d$ are given for $\bm{\mathsf{\Gamma}}$ to satisfy (21), as well as the asymptotic complexity of computing the associated probability distribution, cf. DerezinskiMahoney2021; martinsson2020. We note (by $\dagger$) that the cost of DPP sampling can be reduced, cf. (DerezinskiMahoney2021, Section 5), derezinski2019exact, calandriello2020sampling.

### v2

@@ 3.1 Randomized Matrix Sketching
    where $\bm{\mathsf{D}}\in\mathbb{F}^{m\times m}$ is a random diagonal matrix populated with i.i.d. Rademacher random variables, $\bm{\mathsf{F}}\in\mathbb{F}^{m\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{S}}\in\mathbb{F}^{d\times m}$ is the transpose of a SparseCol matrix with exactly $\zeta$ nonzero entries per column.
These distributions yield randomized DRMs that work well in practice, with theoretical guarantees summarized in Table 1. Gaussian test matrices are OSEs that enjoy a superior lower bound on the requisite embedding dimension $d$ for (21), but the costs of storage and matrix-vector products for Gaussians are much more expensive than for “structured” embeddings like SRTTs or sparse sign matrices. Put simply, Gaussian sketching makes for straightforward analysis, but structured random matrix sketching makes for better performance. The reader is referred to (martinsson2020, Sections 8,9) for details.
@@
Table 1: Summary of coordinate-mixing randomized DRMs commonly used for (row-)sketching rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Theoretical lower bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSE property are provided, as well as the cost to store $\bm{\mathsf{\Gamma}}$ and to apply $\bm{\mathsf{\Gamma}}$ to a vector.
While the embedding dimensions in Table 1 guarantee the OSE property (21), it has been observed that smaller values of $d$ are often sufficient for many linear algebra tasks in practice when utilizing any of the randomized DRMs of Table 1 (with the exception of CountSketch), e.g., $d=k+10$, cf. Sect. 4.1. Recently, this observation was thoroughly investigated from the perspective of oblivious subspace injections, versus embeddings, which relaxes the OSE property as follows. Random matrix $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an oblivious subspace injection (OSI) of $\bm{\mathsf{A}}$ with injectivity $\alpha$ if it satisfies the injectivity property
@@
With the OSI property guiding our choice of random sketching matrix, we enjoy more favorable sketching dimensions for the random matrices considered above. In the OSI framework, Gaussian random matrices are still the gold standard for accuracy, but suffer the same performance drawbacks in terms of storage and matrix-vector product, or mat-vec, costs. Structured random matrices offer significantly better performance, and specifically, sparse random matrices exhibit the best empirical performance, as demonstrated by many practical investigations Chen2025_sparseGPU; ChenakkodDereZinskiDongRudelson2024_OptimalEmbeddingSparse; EpperlyMeierNakatsukasa2024_FastRandomizedLeastSquares; halko2011; KaneNelson2014; martinsson2020; nelson2013; Tropp2015_MatrixConcentration; woodruff2015, and now also supported by the theoretical guarantees of Camano2025_OSI; Tropp2025_OSI in Table 2.
In general, sparse random matrices are recommended for sketching a dense unstructured input matrix, due to the reduced storage costs and significant acceleration afforded by high-performance sparse arithmetic libraries. Compared to the SRTT, the SparseRTT matrix can reduce data movement and better facilitate parallelism at the hardware level. However, the SparseStack matrix is typically faster and thus recommended over SparseRTTs whenever sparse linear algebra packages are available.
As an additional boon, the OSI values of $d$ that are reported in Table 2 are even still too pessimistic in practice. The results of Camano2025_OSI provide compelling evidence that SparseStack matrices are the superior choice of DRM; it is conjectured that SparseStack matrices can achieve the $\frac{1}{2}$-OSI property with $d=O(k)$ and constant sparsity $\zeta=O(1)$. This conjecture is supported numerically, with $d=2k$ and $\zeta=4$ used in the experiments of Camano2025_OSI to demonstrate that the low-rank approximations obtained from the SparseStack sketches are of comparable accuracy to those obtained from Gaussian sketches.
Finally, we note that for input matrices that arise from tensors (cf. Sect. 5.3), we can construct Khatri-Rao sketching matrices that exploit underlying structure Camano2025_OSI, formed as Kronecker products of isotropic random vectors. Often, these random vectors are drawn from (real) Gaussian or Rademacher distributions, but it is advised in Camano2025_OSI that the complex spherical distribution be used to avoid near-zero injectivity in the worst-case. We do not assume any specific isotropic distribution unless otherwise specified.
@@
Table 2: Summary of randomized DRMs for row-sketching rank-$k$ $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Theoretical bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSI property (24) vs. the OSE property (21) are provided, cf. Camano2025_OSI; Tropp2025_OSI. For OSIs, we enforce a constant injectivity parameter, i.e. $\alpha\geq 1/2$, which is generally not attainable for OSEs due to the dilation parameter. For the OSE column of SparseRTTs, the dagger $\dagger$ signifies that a different sparse matrix is used for the OSE property, cf. (Camano2025_OSI, Section 4.1).
@@ 3.2 Randomized Matrix Sampling
In the previous section, we observed that a random sketch has rows (or columns) that are random linear combinations of the rows (or columns) of $\bm{\mathsf{A}}$. However, we can also construct randomized DRMs that sample coordinates of $\bm{\mathsf{A}}$ according to a probability distribution. We use the separate term randomized sampling for these maps.
@@
  • Uniform ((martinsson2020, Section 9.6.2), (tropp2011b, Lemma 3.4), Kannan2017): $p_{i}=1/m.$
  • Squared (row) norms (Frieze2004): $p_{i}=\|\bm{\mathsf{a}}_{i}\|_{2}^{2}/\|\bm{\mathsf{A}}\|^{2}_{\text{F}},$ where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$.
@@
  • Determinantal point processes (DPP) ((DerezinskiMahoney2021, Section 5.2), kulesza2012determinantal): If $S\subseteq[m]$,
@@
    where $|S|=d$, $(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}$ is the sub-matrix of $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\in\mathbb{F}^{m\times m}$ indexed by $S$, and $S$ is selected all at once. However, in practice, rows corresponding to $S$ are selected iteratively, as in Hough_DPP_2005. Let $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}=\sum_{i}\lambda_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ denote the eigendecomposition with $\lambda_{i}\in[0,1]$ for all $i$. Let $s_{i}\sim\textup{Bernoulli}(\lambda_{i})$, and let $\bm{\mathsf{U}}$ be the matrix with columns $\bm{\mathsf{u}}_{i}$. By (Hough_DPP_2005, Theorem 7) and (DerezinskiMahoney2021, Theorem 12), we can use $\sum_{i}s_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ as a proxy for DPP sampling as follows, letting $\bm{\mathsf{v}}_{i}^{*}$ represent the $i$th row of $\bm{\mathsf{U}}$:
@@
We can express the action of randomized sampling as a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$. The lower bounds on the minimal dimension required for a random sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) are summarized in Table 3, along with the cost of computing the associated probabilities. The values of $d$ required for a randomized sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) depend not only on the rank $k$ of $\bm{\mathsf{A}}$ and distortion $\varepsilon$, but also, in some cases, on quantities associated with the matrix $\bm{\mathsf{A}}$ that are usually very expensive to compute.
@@
For uniform sampling, the requisite embedding dimension $d$ depends on the coherence $\mu$ of $\bm{\mathsf{A}}$, defined as $\mu(\bm{\mathsf{A}})=\max_{1\leq i\leq m}\frac{n}{m}l_{i}$, or the maximum leverage score $l_{i}$ in (25). The coherence is a measure of how evenly information in the matrix is distributed; high coherence indicates that its key information is very localized and that uniform sampling may perform poorly as a result; see, e.g., Ipsen14. In spite of this drawback, uniform sampling is a popular choice for large-scale linear algebra computations as a cost-effective OSE. However, it is unclear how to best choose the embedding dimension $d$ in practice since the coherence is typically unavailable, cf. (martinsson2020, Section 9.6.4).
@@
Squared norm sampling and leverage score sampling also yield randomized subspace embeddings, but not oblivious ones, though we can still analyze the embedding dimensions required for them to satisfy (21). For squared norm sampling, the value of $d$ depends on the condition number $\kappa$ of $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$, defined as the ratio of its largest to smallest eigenvalue, which is not readily available in practice. The embedding dimension for leverage score sampling only depends on $k$ and $\epsilon$, but the sampling probabilities are very expensive. In particular, we note that leverage scores are as expensive to directly compute as $\texttt{svd}(\bm{\mathsf{A}})$, rendering their direct computations infeasible for most practical purposes. However, there exist fast methods to approximate them, cf. AlaouiMahoney15; Chen21; DrineasMagdonMahoneyWoodruff12; MuscoMusco17.
@@
DPP sampling has nearly optimal theoretical guarantees, but naive implementations incur a cost of $O(m^{3})$ (after forming $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$). In derezinski2019exact, the DPP-VFX algorithm is presented, which, given access to $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$, samples exactly from a DPP distribution with a pre-processing cost of $O(mk^{10}+k^{15})$. In particular, the DPP-VFX algorithm was the first to achieve $O(m\cdot\textup{poly}(k))$ cost for the first sample and $O(\textup{poly}(k))$ for subsequent samples; however, there is still a runtime bottleneck of $\Omega(m)$ due to the $m$ marginals needed for the sampling distribution. The $\alpha$-DPP algorithm of calandriello2020sampling improves on DPP-VFX by incorporating uniform sub-sampling before computing marginals to avoid the $\Omega(m)$ bottleneck.
@@
Table 3: Summary of randomized sampling distributions commonly used to sample rows of rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Theoretical lower bounds on the embedding dimension $d$ are given for (21), as well as the asymptotic complexity of computing the associated probability distribution, cf. DerezinskiMahoney2021; martinsson2020. We note (by $\dagger$) that the cost of DPP sampling can be reduced, cf. (DerezinskiMahoney2021, Section 5), derezinski2019exact, calandriello2020sampling.